The sequence (x[sub]n[/sub]) is defined by the following formulas for the n[sup]th[/sup] term. Write the first five terms in each case.&
& (x[sub]n[/sub])= ((-1)[sup]n[/sup])/n
1
Expert's answer
2011-09-08T11:43:26-0400
it's not clear enough from the task what x_n is. I suppose x_n=((-1)^n)/notherwise -n/n=-1 and all terms =-1
So x_n=((-1)^n)/n and first five terms(we substitute corresponding
n=1,2,3,4,5 in our formula for x_n)
x_1=-1
x_2=1/2
x_3=-1/3
x_4=1/4
x_5=-1/5